MONT 104Q – Mathematical Journeys: From Known to Unknown Reading and Study Questions on The Adventures of Huckleberry Finn October 2015

1. Why did Twain include the “Notice” and the “Explanatory” on the opening page? Do these give consistent signals to the reader about his (the author’s) intentions regarding the genre of the following book? (That is, can we tell whether this is intended to be a more or less realistic adventure story? a historical novel? a quasi-anthropological investigation of regional dialects? something else?) How would you characterize the genre of this novel?

2. The Adventures of Huckleberry Finn was first published in 1885 and the prefatory note says the action takes place 40 to 50 years before that. What can we say about the historical background to this story and how does that history play a role in the events portrayed?

3. The story is told entirely as a first-person narrative from Huck’s point of view. To what extent do you think we can separate that point of view from Twain’s own point of view? What does the name “Huckleberry Finn” indicate about the main character’s background? How old is he supposed to be in the story? What kind of life has he had up to that point? How would you describe Huck’s personality and character at the start? Can we trust him? Can we accept his version of things, or must we read between his lines? Have you known boys like him? How do our society and school systems tend to treat them today?

4. How are Huck Finn and Tom Sawyer similar? How are they different?

5. How is the runaway slave Jim portrayed?

6. In Jim’s spoken dialogue, it’s clear that Twain was attempting to reproduce the speech patterns of the enslaved African-Americans in Missouri in the 1840’s as precisely as he could. Do you think he was doing this in a malicious way to mock that speech? If so, what underlying attitudes does that reveal? If not, why would he go to that sort of trouble?

7. What do we know about how Twain’s own views about racism and the consequences of slavery in America developed over his lifetime? (To get two “data points” on this, look up the background of his wife, Olivia Langdon, and Warner T. McGuinn, one of the first African-American students admitted to Yale Law School. What is his connection with Twain?)

8. In the course of the story, Huck experiences several crises of conscience leading to decisions to assist Jim (as when Huck tells the two slave hunters that there is “only one” man on the raft and that “He’s white”). Each such decision is more consequential than the previous ones. What are these stages and decisions; when do they occur; and what are their consequences? Does Huck change over the course of the book, and if so, how?

9. Why is Huck’s and Jim’s missing the mouth of the Ohio River in the fog such a big event (chapter XV)? Many commentators about the novel have pointed out the irony that as the novel progresses Huck’s and Jim’s travels take them farther and farther into the deep South and closer to New Orleans. Why is this an ironic feature of the story and how does Twain use that irony?

10. How do the “King” and the “Duke” impact Huck’s and Jim’s life on the raft, their quest for freedom, and the novel’s movement? What are the parallels between the “King’s” and “Duke’s” treatment of Jim in Chapter XXIV and Tom Sawyer’s treatment of him in the final chapters?

11. According to a certain way of understanding the story, Chapter XXXI is the moral climax. What is the key event that happens at that point and what is its meaning? Twain himself said that “Huck Finn is a book of mine where a sound heart and a deformed conscience come into collision and conscience suffers defeat.” What do you think he meant by “a sound heart and a deformed conscience?” How is “conscience” a theme in the novel in general?

12. We know that Twain often recounted the final chapters of Huck Finn, beginning with Huck’s arrival at the Phelps farm, out loud from memory as a sort of humorous “set piece” in the speaking tour he undertook to publicize the book. He says this was always received well by audiences. But do these chapters seem out of place with the rest of the book to you? If we see the end of Chapter XXXI as a climactic moment, do these final chapters build on that climax or lessen its impact?

13. Tom Sawyer’s elaborate charades concerning Jim’s imprisonment and his laughable escape stratagems are borrowed from the genre of historical romance novels like The Count of Monte Cristo and The Man in the Iron Mask. Have you ever read those or seen old movies made from them? If so, do you recognize the similarities? If not, does this whole episode seem humorous to you?

14. One view of the end of Huck Finn might be that the story has a “happy ending.” Huck and Tom have helped win Jim’s freedom and everyone can go on with their lives. But at the very end of the novel, Jim learns a key fact, and then Huck learns another key fact from Jim. What are these facts? How do these revelations create a similarity between Huck’s situation and Jim’s own situation at that point? How does this strike you as a story-telling gambit? Doesn’t it tend to undercut the whole meaning of Jim’s and Huck’s journey? Or is that Twain has a different point here? Has this journey led to any decisive change in Huck’s life?